OpenStudy (anonymous):

HELP PLEASE. For his phone service, Scott pays a monthly fee of 13$, and he pays an additional 0.04 per minute of use. The least he has been charged in a month is 69.24 . What are the possible numbers of minutes he has used his phone in a month? Use m for the number of minutes, and solve your inequality for m .

OpenStudy (anonymous):

here's the equation 69. 24 = 13 + 0.04x 1. the subtraction property of equality, meaning 69. 24 - 13 = 0.04x 2. 56. 24 = 0.04x (x = number of minutes)

OpenStudy (anonymous):

3. just divide 56. 24 by 0.04 and you'll get x =)

OpenStudy (anonymous):

1,406?

OpenStudy (anonymous):

So would it be \[m \ge 1,406\] ?@raimj512

OpenStudy (anonymous):

@raimj512

OpenStudy (anonymous):

is it an inequality? because it would be an exact number. It is 1, 406, so yes, you're right

OpenStudy (anonymous):

yes its an inequality

OpenStudy (anonymous):

I know, I was seeing if you were just guessing, sorry, and yes that's right =)

OpenStudy (anonymous):

lol okay thank you.